OpenStudy (anonymous):

if two painters paint two rooms in two hours then how many painters would be required to paint 18 rooms in six hours? explain..how to use proportionality in this?

OpenStudy (valpey):

Think of it as number of rooms per painter per hour or number of rooms per painter-hour.

OpenStudy (valpey):

In the first instance you have two painters working for two hours so that is 4 "painter-hours". If they can paint two rooms with 4 "painter-hours", then we can infer that each room takes two "painter-hours" to complete.

OpenStudy (valpey):

How many painter-hours does it take to complete 18 rooms?

OpenStudy (anonymous):

i dont understand it please explain again..sorry

OpenStudy (valpey):

1p-h is a measure of work. We know that one room can be completed with 2 p-h's.

OpenStudy (valpey):

1 guy working for 5 hours might get the same amount of work done that five guys working for 1 hour could do. In each case they complete 5 p-h's of work.

OpenStudy (shane_b):

Here's another way to look at it: You have 18 rooms which need to be painted in 6 hours. So you need to paint 3 rooms per hour which is the same as 6 rooms every 2 hours. If 2 painters can paint 2 rooms in 2 hours then how many painters would it take to paint 6 rooms every 2 hours?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

thank you @Shane_B ...thank you@valpey

OpenStudy (valpey):

Cheers!

OpenStudy (shane_b):

Good luck :)

OpenStudy (anonymous):

I got the answer as 54.. Am I right?? Just wanted to know..

OpenStudy (shane_b):

For this question??

OpenStudy (anonymous):

Yes.. I don't know how but I use the proportionality and chain rule for this..

OpenStudy (shane_b):

That answer doesn't even make logical sense :)

OpenStudy (shane_b):

The answer is 6!

OpenStudy (shane_b):

54 painters would paint 27 rooms in the first 2 hours!

OpenStudy (anonymous):

I check it once more.. We have to paint 18 rooms in 6 hours..

OpenStudy (shane_b):

Maybe you've misread the question :)

OpenStudy (anonymous):

I have just solved like this using proportionality....: I show you..

OpenStudy (anonymous):

Yeah I got it, I found the mistake..

OpenStudy (shane_b):

I think you're over-complicating this :)

OpenStudy (anonymous):

Let the painters required be x So, More rooms, More Painters (Direct Proportion) More Hours, Less Painters (Indirect Proportion) So, Rooms Proportion is : 2:18 Hours Proportion is : 6:2 And Painter Proportion is : 2:x So, \[\frac{2}{18} \times \frac{6}{2} = \frac{2}{x}\] \[x=6..\]

OpenStudy (shane_b):

That makes more sense!

OpenStudy (anonymous):

12 ???